A recent study showed that the length of time that juries deliberate on the verdict for civil trials is normally distributed with a mean equal to 15.56 hours with a standard deviation of 3.7 hours. Give this information, what is the probability that a deliberation will last more than 22 hours?


1
Expert's answer
2021-02-28T11:37:43-0500
"p=F(t)"

"t=\frac{22-15.56}{3.7}=1.74"

"p=0.4591"